Tell whether if the following piecewise function is a continuous at a given point or not. (SHOW THE SOLUTION).



1. at x = 4



x + 1 if x < 4


(x - 4)² + 3 if x ≥ 4



2. at x = 0



1/x if x ≤ -1


x² - 2 if x > -1



3. at x = 5



x² - 3 if x ≥ 5


4x + 2 if x < 5




1
Expert's answer
2022-06-08T13:27:10-0400

1.


"\lim\limits_{x\to4^-}f(x)=\lim\limits_{x\to4^-}(x+1)=4+1=5"

"\lim\limits_{x\to4^+}f(x)=\lim\limits_{x\to4^+}((x-4)^2+3)=(4-4)^2+3=3"

"\lim\limits_{x\to4^-}f(x)=5\not=3=\lim\limits_{x\to4^+}f(x)"

"\lim\limits_{x\to4}f(x)=\text{does not exist}"

The function "f(x)" is not continuous at "x=4."

The function "f(x)" has a jump discontinuity at "x=4."


2.


"\lim\limits_{x\to0}f(x)=\lim\limits_{x\to0}(x^2-1)=(0)^2-1=-1=f(0)"

The function "f(x)" is continuous at "x=0."


3.


"\lim\limits_{x\to5^-}f(x)=\lim\limits_{x\to5^-}(4x+2)=4(5)+12=22"

"\lim\limits_{x\to5^+}f(x)=\lim\limits_{x\to5^+}(x^2-3)=(5)^2-3=22"

"\lim\limits_{x\to5^-}f(x)=22=\lim\limits_{x\to5^+}f(x)=>\lim\limits_{x\to5}f(x)=22"

"f(5)=(5)^2-3=22=\lim\limits_{x\to5}f(x)"

The function "f(x)" is continuous at "x=5."